$SPAD/src/input iprntpk.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

)set break resume
)spool iprntpk.output
)set message test on
)set message auto off
)clear all

———-

This package exports but one function. Nonetheless, we had a bug in this code, so a few tests are warranted. The function is iprint, which simply prints a string standard-output without a new line, and flushes the stream.

— * —

--S 1 of 3
)set expose add constructor IPRNTPK
--R   InternalPrintPackage is now explicitly exposed in frame initial
--E 1
--S 2 of 3
iprint("Release the hounds!")
--IRelease the hounds!
--E 2
--S 3 of 3
for i in 1..10 repeat iprint(i::String)
--I12345678910
--E 3
)spool
)lisp (bye)

———-

References^{[1] nothing}

4